# Emmanuel Letellier Université Paris Cité, IMJ-PRG CNRS UMR 7586

### v2

@@ Abstract
In this paper we study the mixed Poincaré polynomials of generic parabolic $\operatorname{PGL}_{2}(\mathbb{C})$-character varieties with coefficients in some local systems arising from the conjugacy class of $\operatorname{PGL}_{2}(\mathbb{C})$ which has a non-connected stabiliser. We give a conjectural formula that we prove to be true under the Euler specialisation. We then prove that these conjectured formulas interpolate the structure coefficients of the two following based rings:
@@ 1 Introduction
For that purpose we need base other than $Cl(G^{F})$ and $\widehat{G^{F}}$ :
@@ 1.1 Review of the case $G=\mathrm{GL}_{n}$
The genericity assumption ensures that the diagonal action of $G$ induces a free action of $\operatorname{PGL}_{n}$. We could have decided to take the quotient by $\operatorname{PGL}_{n}$ instead of $G=\operatorname{GL}_{n}$ but this would require to correct some formulas (like Formula (1.1.2) by a factor $q-1$.
The stack $\mathcal{M}_{\overline{\mathcal{C}}}$, called a $G$-character stack, is the stack of $G$-local systems on the Riemann sphere minus $k$ points $a_{1},\dots,a_{k}$ with local monodromies around $a_{i}$ in $\overline{\mathcal{C}_{i}}$.
@@
$IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)=\sum_{i}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}.$
@@
Then we have [18, Theorem 4.13]
@@
In [18, Conjecture 4.5] we give a conjecture formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ in terms of Macdonald symmetric functions. When the conjugacy classes are semisimple, this was first conjectured in [11] (see §2.2 for a review in the semisimple case). The conjecture is known to be true in some non-trivial examples where we can compute explicitly the weight filtration (see [11, §1.5.3]).
This conjectural formula is proved after the specialisation $t\mapsto-1$ (see [18, Theorem 4.8] and [11, Theorem 1.2.3] in the semisimple case). It is also proved after the specialisation $q\mapsto 1$ which gives the Poincaré series. In the semisimple case this is due to by A. Mellit [27] who followed a strategy used by O. Schiffmann to compute the Poincaré polynomial of the moduli space of semistable Higgs bundles over a smooth projective curve [30]. In the general case, this is due to M. Ballandras [2].
@@
The conjectural formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ implies a conjectural formula for $PIH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q)$ in terms of Hall-Littlewood symmetric functions. In [19, Theorem 6.10.1] we prove that the conjectured formula for the pure part coincides with the inner product
@@
Finally to conclude this part on the $G=\operatorname{GL}_{n}$ case, notice that there is a notion of character-sheaves on the Lie algebra of $G$ due to Lusztig [24] (see also [17]). The Lie algebra analogues $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ of these two rings $\mathcal{A}$ and $\mathcal{B}$ are then isomorphic by an arithmetic Fourier transform (the character-sheaves are precisely the Deligne-Fourier transforms of the intersection cohomology complexes on the Zariski closures of the adjoint orbits of the Lie algebra). The generic structure coefficients of $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ are thus equal. Moreover we can prove that the generic structure coefficients of $\mathcal{B}^{Lie}$ and $\mathcal{B}$ are also equal [19, Theorem 6.9.1] and coincide [11][19] with the Poincaré series (for intersection cohomology) of
@@
where $(\mathcal{O}_{1},\dots,\mathcal{O}_{k})$ runs over the set of generic $k$-tuple of adjoint orbits of ${\rm gl}_{n}(\mathbb{C})$ (unlike the multiplicative case, generic $k$-tuples of adjoint orbits of given Jordan types do not always exist).
Finally, unlike character stacks, the intersection cohomology of $\mathcal{Q}$ is pure [11, Proposition 2.2.6][19, Theorem 7.3.2] and its Poincaré series is conjectured to be the pure part of the mixed Poincaré series of character stacks. This latter conjecture is also a consequence of the conjectural formula for the mixed Poincaré series of generic character stacks.
@@ 1.2 The case where $G=\mathrm{PGL}_{2}$
The aim of this paper is to verify for a group $G\neq G^{*}$ the assumption that the mixed Poincaré series of $G$-character-stack interpolate structure coefficients of the two rings $\mathcal{A}_{G}$ and $\mathcal{B}_{G^{*}}$. We are especially interested in the situation where local monodromies are in conjugacy classes which carry non-trivial $G$-equivariant local systems.
@@
When $K$ is an algebraically closed field of characteristic $\neq 2$, there is only one conjugacy class of $G(K)$ which has a disconnected centraliser. This is the $G(K)$-conjugacy class $\mathcal{C}_{-1}$ of the image in $\operatorname{PGL}_{2}(K)$ of the diagonal matrix with coefficients $1$ and $-1$. There are thus two irreducible (up to isomorphism) $G(K)$-equivariant local systems on $\mathcal{C}_{-1}$ which are naturally parametrised by the set $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{1,\epsilon\}$ of irreducible characters of the group of connected components of the centraliser. We will denote by $\mathcal{L}_{1}$ and $\mathcal{L}_{\epsilon}$ the trivial and non-trivial local systems on $\mathcal{C}_{-1}$ (when the characteristic of $K$ is positive, the local systems considered are in the $\ell$-adic setting).
@@
$\operatorname{PGL}_{2}$-character varieties
@@
Because of the genericity condition this stack is in fact a smooth algebraic variety of dimension $62k-$ 6 (see Theorem 3.2.6).
Choose on each $\mathcal{C}_{i}$ an irreducible $G$-equivariant local system $\mathcal{L}_{i}$ (it is the trivial local system if $\mathcal{C}_{i}$ is non-degenerate and can be either $\mathcal{L}_{1}$ or $\mathcal{L}_{\epsilon}$ if $\mathcal{C}_{i}=\mathcal{C}_{-1}$). This defines a local system $\mathcal{L}_{\mathcal{C}}$ on $\mathcal{M}_{\mathcal{C}}$.
@@
$H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}};q,t)=\sum_{i}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}$
@@
For an $F$-stable $G$-equivariant local system $\mathcal{E}$ on an $F$-stable semisimple conjugacy class $\mathcal{C}$ of $G$, we denote by ${\bf Y}_{(\mathcal{C},\mathcal{E})}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$ the characteristic function of $\mathcal{L}_{\epsilon}$ defined by
@@
(2) If $(\mathcal{C},\mathcal{E})=(\mathcal{C}_{-1},\mathcal{L})$, then
@@
We have a non-canonical bijection
@@
Now the linear characters of $T^{*}{{}^{F}}$ are in bijection with the isomorphism classes of the $F$-stable Kummer local systems on $T^{*}$ (see Proposition 4.3.1).
We thus have a bijection $x\mapsto\mathcal{E}_{x}$ from $\mathbb{F}_{q}^{*}\simeq T^{F}$ to the set of isomorphism classes of $F$-stable Kummer local systems on $T^{*}$.
@@
(2) If $x=-1$, then $\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})$ is an $F$-stable $G^{*}$-equivariant semisimple perverse sheaf whose simple direct factors are naturally paramerised by $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{1,\epsilon\}$ :
$\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})=\mathcal{X}_{1}[3]\oplus\mathcal{X}_{\epsilon}[3].$
The perverse sheaves $\mathcal{X}_{1}[3]$ and $\mathcal{X}_{\epsilon}[3]$ are other examples of character-sheaves. However unlike the ones above their characteristic functions are not irreducible characters (see the end of §4.5).
@@
$\Phi:\left\{(\mathcal{C}_{x},\overline{{\mathbb{Q}}}_{\ell})\,|\,x\neq 1,-1\right\}\sqcup\{(\mathcal{C}_{-1},\mathcal{L}_{r})\,|\,r=1,\epsilon\}\simeq\left\{\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})\,|\,x\neq 1,-1\right\}\sqcup\left\{\mathcal{X}_{r}[3]\,|\,r=1,\epsilon\right\}$
where $\mathcal{C}_{x}$ is the (semisimple) $G$-conjugacy class of $x\in\mathbb{F}_{q}^{*}\simeq T^{F}$.
@@
Choose a generic $k$-tuple of $F$-stable semisimple regular conjugacy classes of $G(\overline{\mathbb{F}}_{q})$ whose $i$-th coordinate is degenerate (i.e. equals to $\mathcal{C}_{-1}$) if and only if $\mathcal{C}_{i}$ is degenerate. To ease the notation we will denote this $k$-tuple again by $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$. On each conjugacy class we choose the same local system as in the complex case and denote it again by $\mathcal{L}_{i}$. We then denote by ${\bf Y}_{i}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}\simeq\mathbb{C}$ the characteristic function ${\bf Y}_{(\mathcal{C}_{i},\mathcal{L}_{i})}$ and by ${\bf X}_{i}$ is the characteristic function of the character-sheaf of $G^{*}$ which corresponds to $(\mathcal{C}_{i},\mathcal{L}_{i})$ under the bijection $\Phi$.
@@
(2) The conjectured formula for the pure part $PH_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{L}_{\mathcal{C}};q)$ equals
@@ 2.1.1 Polyomial count varieties
In this case, we have the following [20, Theorem 2.2.6].
@@
Indeed, let $f:X_{R}\to\operatorname{Spec}(R)$ be the structural morphism. The complex $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constructible, see for instance [5, Chapter 2]. In particular, there exists a non-empty open $U\subseteq\operatorname{Spec}(R)$ such that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant. Denote by $\xi:\operatorname{Spec}(\mathbb{C})\to\operatorname{Spec}(R)$ the (geometric) generic point of $\operatorname{Spec}(R)$ coming from the embedding $R\subseteq\mathbb{C}$ and, for any $\varphi:R\to\mathbb{F}_{q}$, denote by $\xi_{\varphi}:\operatorname{Spec}(\overline{\mathbb{F}}_{q})\to\operatorname{Spec}(R)$ the corresponding geometric point.
If $\operatorname{Im}(\xi_{\varphi})\in U$, from the fact that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant and from proper base change, we have the following chain of isomorphisms:
@@
The results of [6, Theorem 14] show that the isomorphism (2.1.2) preserves the weight filtration on both sides. In particular, we have an equality

### v3

@@ Abstract
In this paper we study the mixed Poincaré polynomials of generic $\operatorname{PGL}_{2}(\mathbb{C})$-character stacks with coefficients in some local systems arising from the conjugacy class of $\operatorname{PGL}_{2}(\mathbb{C})$ which has a non-connected stabiliser. We give a conjectural formula that we prove to be true under the Euler specialisation. We then prove that these conjectured formulas interpolate the structure coefficients of the two following based rings:
@@ 1 Introduction
For that purpose we need a base other than $Cl(G^{F})$ and $\widehat{G^{F}}$ :
@@ 1.1 Review of the case $G=\mathrm{GL}_{n}$
The genericity assumption ensures that the diagonal action of $G$ induces a free action of $\operatorname{PGL}_{n}$. In particular, the stack $\mathcal{M}_{\overline{\mathcal{C}}}$ is a $\mathbb{G}_{m}$-gerbe over the smooth variety $M_{\overline{\mathcal{C}}}$ obtained by taking the quotient by $\operatorname{PGL}_{n}$. We still take the quotient by $G=\operatorname{GL}_{n}$ in the definition since otherwise we would have to to correct some formulas (like Formula (1.1.2) by a factor $q-1$.
The stack $\mathcal{M}_{\overline{\mathcal{C}}}$, called a $G$-character stack, is the moduli stack of $G$-local systems on the Riemann sphere $\mathbb{P}^{1}_{\mathbb{C}}$ minus $k$ points $a_{1},\dots,a_{k}$ with local monodromies around $a_{i}$ in $\overline{\mathcal{C}_{i}}$.
@@
$IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)=\sum_{i,r}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}.$
@@
Then we have [17, Theorem 4.13]
@@
In [17, Conjecture 4.5] we give a conjecture formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ in terms of Macdonald symmetric functions. When the conjugacy classes are semisimple, this was first conjectured in [11] (see §2.2 for a review in the semisimple case). The conjecture is known to be true in some non-trivial examples where we can compute explicitly the weight filtration (see [11, §1.5.3]).
This conjectural formula is proved after the specialisation $t\mapsto-1$ (see [17, Theorem 4.8] and [11, Theorem 1.2.3] in the semisimple case). It is also proved after the specialisation $q\mapsto 1$ which gives the Poincaré series. In the semisimple case this is due to by A. Mellit [26] who followed a strategy used by O. Schiffmann to compute the Poincaré polynomial of the moduli space of semistable Higgs bundles over a smooth projective curve [29]. In the general case, this is due to M. Ballandras [2].
@@
The conjectural formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ implies a conjectural formula for $PIH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q)$ in terms of Hall-Littlewood symmetric functions. In [18, Theorem 6.10.1] we prove that the conjectured formula for the pure part coincides with the inner product
@@
Finally to conclude this part on the $G=\operatorname{GL}_{n}$ case, notice that there is a notion of character-sheaves on the Lie algebra of $G$ due to Lusztig [23] (see also [16]). The Lie algebra analogues $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ of these two rings $\mathcal{A}$ and $\mathcal{B}$ are then isomorphic by an arithmetic Fourier transform (the character-sheaves are precisely the Deligne-Fourier transforms of the intersection cohomology complexes on the Zariski closures of the adjoint orbits of the Lie algebra). The generic structure coefficients of $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ are thus equal. Moreover we can prove that the generic structure coefficients of $\mathcal{B}^{Lie}$ and $\mathcal{B}$ are also equal [18, Theorem 6.9.1] and coincide [11][18] with the Poincaré series (for intersection cohomology) of
@@
where $(\mathcal{O}_{1},\dots,\mathcal{O}_{k})$ runs over the set of generic $k$-tuple of adjoint orbits of $\mathfrak{gl}_{n}(\mathbb{C})$ (unlike the multiplicative case, generic $k$-tuples of adjoint orbits of given Jordan types do not always exist).
Finally, unlike character stacks, the intersection cohomology of $\mathcal{Q}$ is pure [11, Proposition 2.2.6][18, Theorem 7.3.2] and its Poincaré series is conjectured to be the pure part of the mixed Poincaré series of character stacks. This latter conjecture is also a consequence of the conjectural formula for the mixed Poincaré series of generic character stacks.
@@ 1.2 The case where $G=\mathrm{PGL}_{2}$
The aim of this paper is to verify for a group $G\neq G^{*}$ the assumption that the mixed Poincaré series of $G$-character stacks interpolate structure coefficients of the two rings $\mathcal{A}_{G}$ and $\mathcal{B}_{G^{*}}$. We are especially interested in the situation where local monodromies are in conjugacy classes which carry non-trivial $G$-equivariant local systems.
@@
When $K$ is an algebraically closed field of characteristic $\neq 2$, there is only one conjugacy class of $G(K)$ which has a disconnected centraliser. This is the $G(K)$-conjugacy class $\mathcal{C}_{-1}$ of the image in $\operatorname{PGL}_{2}(K)$ of the diagonal matrix with coefficients $1$ and $-1$. There are thus two irreducible (up to isomorphism) $G(K)$-equivariant local systems on $\mathcal{C}_{-1}$ which are naturally parametrised by the set $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{{\rm Id},\epsilon\}$ of irreducible characters of the group of connected components of the centraliser. We will denote by $\mathcal{L}_{\epsilon}$ the non-trivial equivariant irreducible local systems on $\mathcal{C}_{-1}$ (when the characteristic of $K$ is positive, the local systems considered are in the $\ell$-adic setting).
@@
$\operatorname{PGL}_{2}$-character stacks
@@
Choose on each $\mathcal{C}_{i}$ an irreducible $G$-equivariant local system $\mathcal{L}_{i}$ (it is the trivial local system if $\mathcal{C}_{i}$ is non-degenerate and can be either the trivial local system or $\mathcal{L}_{\epsilon}$ if $\mathcal{C}_{i}=\mathcal{C}_{-1}$). This defines a local system $\mathcal{L}_{\mathcal{C}}$ on $\mathcal{M}_{\mathcal{C}}$.
@@
$H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}};q,t)=\sum_{i,r}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}$
@@
For an $F$-stable $G$-equivariant local system $\mathcal{E}$ on an $F$-stable semisimple conjugacy class $\mathcal{C}$ of $G$, we denote by ${\bf Y}_{(\mathcal{C},\mathcal{E})}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$ the characteristic function of $\mathcal{E}$ defined by
@@
(2) If $(\mathcal{C},\mathcal{E})=(\mathcal{C}_{-1},\mathcal{L}_{\epsilon})$, then
@@
We have a non-canonical group isomorphism
@@
Now the group of linear characters of $T^{*}{{}^{F}}$ is isomorphic with the group of isomorphism classes of the $F$-stable Kummer local systems on $T^{*}$ (see Proposition 4.3.1).
We thus have an isomorphism $x\mapsto\mathcal{E}_{x}$ from $\mathbb{F}_{q}^{*}\simeq T^{F}$ to the group of isomorphism classes of $F$-stable Kummer local systems on $T^{*}$.
@@
(2) If $x=-1$, then $\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})$ is an $F$-stable $G^{*}$-equivariant semisimple perverse sheaf whose simple direct factors are naturally parametrised by $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{{\rm Id},\epsilon\}$ :
$\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})=\mathcal{X}_{{\rm Id}}[3]\oplus\mathcal{X}_{\epsilon}[3].$
The perverse sheaves $\mathcal{X}_{{\rm Id}}[3]$ and $\mathcal{X}_{\epsilon}[3]$ are other examples of character-sheaves. However unlike the ones above their characteristic functions are not irreducible characters (see the end of §4.5).
@@
$\Phi:\left\{(\mathcal{C}_{x},\overline{{\mathbb{Q}}}_{\ell})\,|\,x\neq 1,-1\right\}\sqcup\{(\mathcal{C}_{-1},\mathcal{L}_{r})\,|\,r={\rm Id},\epsilon\}\simeq\left\{\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})\,|\,x\neq 1,-1\right\}\sqcup\left\{\mathcal{X}_{r}[3]\,|\,r={\rm Id},\epsilon\right\}$
where $\mathcal{C}_{x}$ is the (semisimple) $G$-conjugacy class of $x\in\mathbb{F}_{q}^{*}\simeq T^{F}$, and where $\mathcal{L}_{{\rm Id}}$ is the constant sheaf on $\mathcal{C}_{-1}$.
@@
Choose a generic $k$-tuple of $F$-stable semisimple regular conjugacy classes of $G(\overline{\mathbb{F}}_{q})$ whose $i$-th coordinate is degenerate (i.e. equals to $\mathcal{C}_{-1}$) if and only if $\mathcal{C}_{i}$ is degenerate. To ease the notation we will denote this $k$-tuple again by $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$. On each conjugacy class we choose the same local system as in the complex case and denote it again by $\mathcal{L}_{i}$. We then denote by ${\bf Y}_{i}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}\simeq\mathbb{C}$ the characteristic function ${\bf Y}_{(\mathcal{C}_{i},\mathcal{L}_{i})}$ and by ${\bf X}_{i}$ the characteristic function of the character-sheaf of $G^{*}$ which corresponds to $(\mathcal{C}_{i},\mathcal{L}_{i})$ under the isomorphism $\Phi$.
@@
(2) The conjectured formula for the pure part
$PH_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{L}_{\mathcal{C}};q):=\sum_{i}{\rm dim}\,\left(W^{i}_{i}/W^{i}_{i-1}\right)\,q^{i/2}$
equals
@@ 2.1 Preliminaries on cohomology
In this case, we have the following [19, Theorem 2.2.6].
@@
Indeed, let $f:X_{R}\to\operatorname{Spec}(R)$ be the structural morphism. The complex $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constructible, see for instance [4, Chapter 2]. In particular, there exists a non-empty open $U\subseteq\operatorname{Spec}(R)$ such that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant. Denote by $\xi:\operatorname{Spec}(\mathbb{C})\to\operatorname{Spec}(R)$ the (geometric) generic point of $\operatorname{Spec}(R)$ coming from the embedding $R\subseteq\mathbb{C}$ and, for any $\varphi:R\to\mathbb{F}_{q}$, denote by $\xi_{\varphi}:\operatorname{Spec}(\overline{\mathbb{F}}_{q})\to\operatorname{Spec}(R)$ the corresponding geometric point.
If $\operatorname{Im}(\xi_{\varphi})\in U$, from the fact that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant and from the proper base change theorem, we have the following chain of isomorphisms:
@@
The results of [5, Theorem 14] show that the isomorphism (2.1.2) preserves the weight filtration on both sides. In particular, we have an equality